# Prioritized-MVBA: A New Approach to Design an Optimal Asynchronous Byzantine Agreement Protocol

### v1

@@ Abstract
The multi-valued byzantine agreement protocol (MVBA) in the authenticated setting has been widely used as a core to design atomic broadcast and fault-tolerant state machine replication protocols in asynchronous networks. Originating from the seminal work of Cachin et al. [13], subsequent research endeavors have sought to optimize protocol efficiency in terms of communication complexity. Notable advancements following Cachin’s contributions include: i) VABA [23], requiring multiple protocol instances to achieve agreement on a party’s request, and ii) Dumbo-MVBA [29], employing a cryptographic asynchronous dispersal and recovery methods to manage communication complexity alongside additional computational and communication rounds overheads.
Our objective is to devise an MVBA protocol that achieves agreement in each instance without extra computation and communication rounds while maintaining the optimal metrics. Through meticulous research, we analyze the network/adversary’s message distribution patterns and observe how a party’s message reception varies with the number of request senders. We ascertain that opting for a subset of parties ($f+1$, ensuring the inclusion of at least one honest party, at most $f$ number of parties can be faulty ) rather than the entire set ($n$) can facilitate the design of a protocol meeting the aforementioned criteria.
Central to our design approach is the introduction of the committee in the classic MVBA protocol, wherein a randomly selected subset of ($f+1$, where $n=3f+1$) parties get selected and simultaneously broadcast their requests (transactions) to gather verifiable proofs. Successive distributions of these proofs afford us the necessary properties to employ the asynchronous binary Byzantine agreement (ABBA) protocol for reaching an agreement on a selected party’s requests. By integrating the committee and ABBA protocols, we devise the optimal MVBA protocol, termed pMVBA (Prioritized-MVBA). This protocol exhibits resilience to tolerate up to $\lfloor\frac{n}{3}\rfloor$ Byzantine failures, with an expected runtime of $O(1)$, optimal message complexity of $O(n^{2})$, and optimal communication complexity $O((l+\lambda)n^{2})$ (where $n$ represents the number of parties, $l$ denotes the bit length of the input, and $\lambda$ signifies the security parameter).
@@ 1 Introduction
Byzantine Agreement (BA) protocols are fundamental to implementing decentralized infrastructures, particularly in the context of blockchain and decentralized applications. These protocols ensure agreement among parties, even in the presence of malicious actors, making them crucial for maintaining the integrity and functionality of distributed systems. The renewed interest in BA protocols has been driven by the success of Bitcoin [35] and other decentralized applications [44], which rely heavily on these protocols for achieving agreement [14]. Despite the existence of practical BA protocols [1, 33, 5, 37, 41, 31], their reliance on time parameters for ensuring liveness poses challenges in designing agreement protocols. The FLP impossibility result [32] demonstrates that deterministic BA protocols cannot guarantee agreement in asynchronous networks. This necessitates the development of efficient asynchronous Byzantine Agreement protocols [23, 42, 11, 5, 29, 10, 39, 45, 22] that can overcome the limitations of existing approaches, particularly in terms of communication complexity.
Achieving agreement in asynchronous networks efficiently remains a significant challenge. Classic Multi-Valued Byzantine Agreement (MVBA) protocol like Cachin’s [13], while effective, suffer from high communication complexity, often involving terms like $O(n^{3})$. This high complexity is impractical for large-scale systems, necessitating new approaches to reduce communication overhead while ensuring optimal resilience and correctness. Cachin et al. [13] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision. Cachin’s MVBA protocol provides polynomial-time complexity but suffers from high communication overhead, with a complexity of $O(n^{3})$. VABA [23] introduced a view-based approach to eliminate the $O(n^{3})$ term, while Dumbo-MVBA [29] used erasure codes to handle large input sizes more efficiently, reducing the communication complexity to $O(ln+\lambda n^{2})$. However, these optimizations often introduced additional rounds of communication and computation.
@@
The main contributions of this paper are as follows:
  • We introduce a novel committee selection protocol that ensures the presence of at least one honest party in the selected subset, enhancing the protocol’s resilience against Byzantine failures.
  • We propose the pMVBA protocol, which reduces the communication complexity from $O(n^{3})$ to $O((l+\lambda)n^{2})$ by dynamically selecting a subset of parties to broadcast their proposals.
  • Our protocol integrates the ABBA protocol to achieve agreement with probability $1$ on the proposals broadcast by committee members, ensuring robust agreement mechanisms.
  • We provide a thorough theoretical analysis demonstrating that the pMVBA protocol achieves an expected constant asynchronous round count while maintaining optimal resilience and correctness.
  • We present extensive case studies and performance evaluations, showing that the pMVBA protocol is effective and scalable for large-scale decentralized applications, outperforming traditional MVBA protocols in terms of communication overhead and efficiency.
@@ 2 Related Work
Byzantine Agreement Protocols. Byzantine Agreement (BA) protocols have long been a cornerstone of fault-tolerant distributed computing. The Byzantine Generals Problem, introduced by Lamport et al. [26], laid the theoretical foundation for BA protocols, emphasizing the challenges of achieving an agreement in the presence of faulty or malicious nodes. Early BA protocols, such as those by Pease et al. [38], focused on synchronous networks where communication occurs in fixed rounds. However, these protocols often relied on time assumptions for ensuring liveness.
Practical Byzantine Fault Tolerance (PBFT), proposed by Castro and Liskov [15], was a significant advancement in making BA protocols practical for real-world applications. PBFT introduced an efficient consensus protocol for partially-synchronous networks, ensuring safety and liveness even with up to $f<\frac{n}{3}$ Byzantine faults. Several modern BA protocols build upon the principles of PBFT, including HotStuff [46] and Tendermint [12], optimizing various aspects such as communication complexity and responsiveness.
The FLP impossibility result [18] demonstrated that no deterministic BA protocol could guarantee an agreement in an asynchronous network with a single faulty node. This led to the exploration of probabilistic and randomized approaches for ABA. Ben-Or’s protocol [8] was one of the earliest randomized BA protocols, using coin-flipping to break symmetry among nodes. Recent protocols like HoneyBadgerBFT [34] and Dumbo [28] have focused on optimizing ABA for high throughput and low latency.
Multi-Valued Byzantine Agreement Protocols. Cachin et al. [13] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision. Cachin’s MVBA protocol provides polynomial-time complexity but suffers from high communication overhead, with a complexity of $O(ln^{2}+\lambda n^{2}+n^{3})$.
Subsequent works, such as VABA [23] and Dumbo-MVBA [29], aimed to reduce the communication complexity of MVBA. VABA introduced a view-based approach to eliminate the $O(n^{3})$ term, while Dumbo-MVBA used erasure codes to handle large input sizes more efficiently, reducing the communication complexity to $O(ln+\lambda n^{2})$. However, these optimizations often introduced additional rounds of communication and expensive cryptograhic computations.
@@
Committee-Based Protocols. Committee-based protocols have emerged as a promising approach to reduce communication complexity in BA protocols. Algorand [20] introduces the concept of committee members, wherein a predefined set of parties collaborates to reach an agreement. Algorand operates within a permission-less (non-trusted) setting. In a similar vein, COINcidence [19] presents a protocol within a trusted setup, leveraging committee-based participation to achieve agreement with high probability.
Our approach diverges from both Algorand and COINcidence. Unlike Algorand, we operate within a trusted setup, and unlike both Algorand and COINcidence, our protocol involves selected parties undertaking a singular step. Furthermore, our protocol guarantees an agreement on a single party’s requests with a probability of 1. Conversely, FasterDumbo [11] utilizes the committee concept to minimize the number of components in the agreement process, yielding outputs from $n-f$ parties’ requests. Beat [42] amalgamates the most effective elements of HoneyBadgerBFT’s concrete implementation to showcase superior performance across varied settings.
Optimistic Fastlane Mechanisms. Certain works, such as those by Victor et al. [25] and Cachin et al. [17], have explored the incorporation of an optimistic ’fastlane’ mechanism, designed for operation within partially synchronous networks. However, in instances of sluggish agreement processes, these mechanisms transition to asynchronous Byzantine agreement protocols. Recent endeavors, such as those by Joleton [39] and Boltdumbo [45], have adopted a similar strategy. Our work remains pertinent within this context as we rely on asynchronous protocols when the network exhibits Byzantine behavior, particularly within the pessimistic path.
Given our protocol’s focus on outputting requests from a single party, we omit comparisons with DAG-rider [22] and Aleph [21], which utilize directed acyclic graphs for agreement alongside sequential ACS. Additionally, recent efforts have aimed to eliminate the private setup phase in asynchronous Byzantine agreement protocols, emphasizing asynchronous distributed key generation [40, 30, 7].
@@ 3 System Model
We use the standard notion [16, 13] to describe a distributed algorithm involving $n$ parties and an adversary in an authenticated setting. This section explains the foundational assumptions, network model, adversary model, computational model, and cryptographic tools used in our protocol.
@@ 3.1 System and Network Assumptions
Parties and Their Setup: There are $n$ parties, denoted as $p_{1},p_{2},\ldots,p_{n}$. We consider a trusted setup of threshold cryptosystems. The system provides a secret key and a public key to each party before the protocol starts. A party uses its secret key to sign a message, and other parties use the corresponding public key to verify the signed message. The generation and distribution of these keys are out of the scope of this paper. We follow standard literature for key generation and distribution and refer interested readers to [9, 4, 6]. We use the node and party alternatively throughout the paper.
@@
Computational Model: We adopt standard modern cryptographic assumptions and definitions from [16, 13]. The assumptions allow the parties and the adversary to be probabilistic polynomial-time interactive Turing machines. This means that upon receiving a message, a party carries out some computations, changes its state, generates outgoing messages if required, and waits for the next incoming message. To rule out infinite protocol executions and restrict the run time of the adversary, we require the message bits generated by honest parties to be probabilistic uniformly bounded by a polynomial in the security parameter $\lambda$. Therefore, we assume that the number of parties $n$ is bounded by a polynomial in $\lambda$.
@@ 3.2 Validated Asynchronous Byzantine Agreement:
Multi-valued Byzantine Agreement (MVBA) allows parties to agree on an arbitrary string $\{0,1\}^{l}$, which must satisfy a predefined validity condition before the honest parties agree on it. This protocol ensures that the agreed-upon values are valid even when inputs come from malicious parties. This property is crucial for ensuring the protocol’s correctness. Therefore, it is called the validated asynchronous byzantine agreement [13, 24, 47].
@@ 3.3 Preliminaries
Guo et al. [11] constructed the $(1,\kappa,\epsilon)$-Committee Selection protocol using a threshold coin-tossing mechanism (see Appendix A.3), which is derived from threshold signatures. The protocol ensures that at least one honest party is a committee member with overwhelming probability $1-\epsilon-\text{neg}(\lambda)$, where $\text{neg}(\lambda)$ is a negligible function in the cryptographic security parameter $\lambda$, and $\epsilon$ is $exp(-\Omega\kappa)$.
Cryptographic Abstractions: We design a distributed algorithm in authenticated settings where we use robust, non-interactive threshold signatures to authenticate messages and a threshold coin-tossing protocol to select parties randomly [16].
@@
    Threshold Signature Scheme: We utilize a threshold signature scheme introduced in [43, 16]. The basic idea is that there are $n$ parties, up to $f$ of which may be corrupted. The parties hold shares of the secret key of a signature scheme and may generate shares of signatures on individual messages. $t$ signature shares are both necessary and sufficient to construct a signature where $f<t\leq(n-f)$. The threshold signature scheme also provides a public key $pk$ along with secret key shares $sk_{1},\ldots,sk_{n}$, a global verification key $vk$, and local verification keys $vk_{1},\ldots,vk_{n}$. Initially, a party $p_{i}$ has information on the public key $vk$, a secret key $sk_{i}$, and the verification keys for all the parties’ secret keys. We describe the security properties of the scheme and related algorithms in Appendix A.2.
@@
    Threshold Coin-Tossing Scheme: The threshold coin-tossing scheme, as introduced in [43, 16], involves parties holding shares of a pseudorandom function $F$ that maps the name $C$ (an arbitrary bit string) of a coin. We use a distributed pseudorandom function as a coin that produces $k^{\prime\prime}$ random bits simultaneously. The name $C$ is necessary and sufficient to construct the value $F(C)\in\{0,1\}^{k^{\prime\prime}}$ of the particular coin. The parties may generate shares of a coin — $t$ coin shares are both necessary and sufficient where $f<t\leq n-f$, similar to threshold signatures. The generation and verification of coin-shares are also non-interactive. We describe the security properties of the scheme and related algorithms in Appendix A.3.
4 The Proposed Protocol
(a) An overview of the proposed protocol framework, highlighting our contributions: committee selection, the VCBC protocol for committee members, recommendations, and Sequential ABBA.

### v2

@@ Abstract
Multi-valued Byzantine agreement (MVBA) protocols are critical components in designing atomic broadcast and fault-tolerant state machine replication protocols in asynchronous networks. While these protocols have seen significant advancements, challenges remain in optimizing their communication and computation efficiency without sacrificing performance. In this paper, we address the challenge of achieving agreement in MVBA without incurring extra computation and communication rounds. Our approach leverages an analysis of message distribution patterns in asynchronous networks, observing that a subset of $f+1$ parties (where at least one party is honest), can achieve an agreement more efficiently than relying on all $n$ parties, where $n=3f+1$, $f$ is the maximum number of faulty parties. We introduce a novel protocol, Prioritized-MVBA (pMVBA), which integrates a committee-based selection process and the asynchronous binary Byzantine agreement (ABBA) protocol. In this design, a randomly selected subset of $f+1$ parties broadcast their requests, collect verifiable proofs, and utilize these proofs within the ABBA framework to reach an agreement. The proposed pMVBA protocol is resilient to up to $\lfloor\frac{n}{3}\rfloor$ Byzantine failures and achieves optimal performance, with an expected runtime of $O(1)$, message complexity of $O(n^{2})$, and communication complexity of $O((l+\lambda)n^{2})$ (similar to state-of-the-art protocols), where $n$ is the number of parties, $l$ is the input bit length, and $\lambda$ is the security parameter. The protocol attains $68\%$ higher throughput and $18\%$ lower latency than the classic MVBA protocol (the most efficient protocol).
@@ 1 Introduction
Byzantine Agreement (BA) protocols are fundamental to implementing decentralized infrastructures, particularly in the context of blockchain and decentralized applications. These protocols ensure agreement among parties, even in the presence of malicious actors, making them crucial for maintaining the integrity and functionality of distributed systems. The renewed interest in BA protocols has been driven by the success of Bitcoin [SATOSHI08] and other decentralized applications [SC], which rely heavily on these protocols for achieving agreement [CACHIN17]. Despite the existence of practical BA protocols [ABRAHAM20, BYZ08, HONEYBADGER01, PASS17, SHRESTHA20, Hotstuff01], their reliance on time parameters for ensuring liveness poses challenges in designing agreement protocols. The FLP impossibility result [CONS03] demonstrates that deterministic BA protocols cannot guarantee agreement in asynchronous networks. This necessitates the development of efficient asynchronous Byzantine Agreement protocols [BYZ17, DUAN18, FASTERDUMBO, HONEYBADGER01, LU20, SPEEDINGDUMBO, Joleton, BOLTDUMBO, DAG] that can overcome the limitations of existing approaches, particularly in terms of communication complexity.
Achieving agreement in asynchronous networks efficiently remains a significant challenge. Classic Multi-Valued Byzantine Agreement (MVBA) protocol like Cachin’s [CACHIN01], while effective, suffer from high communication complexity, often involving terms like $O(n^{3})$. This high complexity is impractical for large-scale systems, necessitating new approaches to reduce communication overhead while ensuring optimal resilience and correctness. Cachin et al. [CACHIN01] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision. Cachin’s MVBA protocol provides polynomial-time complexity but suffers from high communication overhead, with a complexity of $O(n^{3})$. VABA [BYZ17] introduced a view-based approach to eliminate the $O(n^{3})$ term, while Dumbo-MVBA [LU20] used erasure codes to handle large input sizes more efficiently, reducing the communication complexity to $O(ln+\lambda n^{2})$. However, these optimizations often introduced additional rounds of communication and computation.
@@
(a) O(n) bottleneck
(b) No bottleneck
Figure 1: Comparison of two protocols. We show the time complexity of each sub-protocol and the step where our protocol wins in terms of communication. The txtx represents the transaction, the tsts represents the threshold signature. $1,1,1,0$ represents the bit array and each color for the particular party and its messages.
The main contributions of this paper are as follows: The main contributions of this project are as follows:
  • We propose the pMVBA protocol, which reduces the communication complexity from $O(n^{3})$ [SECURE03] to $O((l+\lambda)n^{2})$ by dynamically selecting a subset of parties using the cryptographic coin-tossing technique to broadcast their proposals.
  • We introduce a committee in the classical MVBA protocol [SECURE03] that ensures that the protocol reaches an agreement with probability $1$, unlike other committee-based protocols. We replace the $O(n)$ bit Commit step of the classic protocol with $O(1)$ Recommend step.
  • We integrate the ABBA biased towards $1$ protocol to mitigate the effect of the reduction in the total number of proposals due to the introduction of the committee.
  • We present proof, extensive case studies, and experimental results to show that our approach (committee and Recommend step) fills the requirement of the ABBA protocol to reach an agreement on a committee member’s proposal.
  • We implement pMVBA and extensively test it among $n=4$, 77 or $10$ nodes in gRPC based simulator, making detailed Comparison to the state-of-the-art MVBA protocols, including ClassicMVBA[SECURE03] (CKSP01), and one recent result, VABA (Abraham et al. [BYZ17]). At all system scales, the peak throughput of pMVBA is better than any of the other tested MVBA protocols (up to $180\%$). (see Table 1).
    More importantly, the latency of pMVBA protocol is significantly less than that of others (e. g. VABA [BYZ17], CKSP01 [SECURE03]). Actually, the latency of pMVBA is nearly independent of its throughput, indicating the effectiveness of the protocol to resolve the throughput-latency tension lying in the previous protocols. As shown in Figure 2, we see that pMVBA has the lowest increment in terms of latency, and the increment ratio is also lowest among other protocols. This result indicates the effectiveness of the proposed MVBA protocol.
    Table 1: In comparison with existing MVBA protocols.
    Protocols Peak throughput (tps) Latency at peak-tps(sec)
    n=4 n=7 n = 10 n=4 n=7 n = 10
    VABA [BYZ17] 30.5 15.1 9 16.39 33 55.35
    CKSP01 [SECURE03] 53.47 23.8 12 9.35 21 41.92
    pMVBA (ours) 85.67 37.87 22 5.84 13.2 22.89
    Figure 2: Latency increment of MVBA protocols when throughput increases from minimum to maximum (n=10)
@@ 2 Related Work
Byzantine Agreement Protocols. Byzantine Agreement (BA) protocols have long been a cornerstone of fault-tolerant distributed computing. The Byzantine Generals Problem, introduced by Lamport et al. [LAMPORT82], laid the theoretical foundation for BA protocols, emphasizing the challenges of achieving an agreement in the presence of faulty or malicious nodes. Early BA protocols, such as those by Pease et al. [PEASE80], focused on synchronous networks where communication occurs in fixed rounds. However, these protocols often relied on time assumptions for ensuring liveness.
Practical Byzantine Fault Tolerance (PBFT), proposed by Castro and Liskov [CASTRO99], was a significant advancement in making BA protocols practical for real-world applications. PBFT introduced an efficient consensus protocol for partially-synchronous networks, ensuring safety and liveness even with up to $f<\frac{n}{3}$ Byzantine faults. Several modern BA protocols build upon the principles of PBFT, including HotStuff [HOTSTUFF19] and Tendermint [TENDERMINT14], optimizing various aspects such as communication complexity and responsiveness.
The FLP impossibility result [FLP85] demonstrated that no deterministic BA protocol could guarantee an agreement in an asynchronous network with a single faulty node. This led to the exploration of probabilistic and randomized approaches for ABA. Ben-Or’s protocol [BENOR83] was one of the earliest randomized BA protocols, using coin-flipping to break symmetry among nodes. Recent protocols like HoneyBadgerBFT [MILLER16] and Dumbo [DUMBO20] have focused on optimizing ABA for high throughput and low latency.
Multi-Valued Byzantine Agreement Protocols. Cachin et al. [CACHIN01] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision. Cachin’s MVBA protocol provides polynomial-time complexity but suffers from high communication overhead, with a complexity of $O(ln^{2}+\lambda n^{2}+n^{3})$.
Subsequent works, such as VABA [BYZ17] and Dumbo-MVBA [LU20], aimed to reduce the communication complexity of MVBA. VABA introduced a view-based approach to eliminate the $O(n^{3})$ term, while Dumbo-MVBA used erasure codes to handle large input sizes more efficiently, reducing the communication complexity to $O(ln+\lambda n^{2})$. However, these optimizations often introduced additional rounds of communication and expensive cryptograhic computations.
@@
Committee-Based Protocols. Committee-based protocols have emerged as a promising approach to reduce communication complexity in BA protocols. Algorand [BYZ21] introduces the concept of committee members, wherein a predefined set of parties collaborates to reach an agreement. Algorand operates within a permission-less (non-trusted) setting. In a similar vein, COINcidence [COINCIDENCE] presents a protocol within a trusted setup, leveraging committee-based participation to achieve agreement with high probability.
Our approach diverges from both Algorand and COINcidence. Unlike Algorand, we operate within a trusted setup, and unlike both Algorand and COINcidence, our protocol involves selected parties undertaking a singular step. Furthermore, our protocol guarantees an agreement on a single party’s requests with a probability of 1. Conversely, FasterDumbo [FASTERDUMBO] utilizes the committee concept to minimize the number of components in the agreement process, yielding outputs from nfn-f parties’ requests. Beat [DUAN18] amalgamates the most effective elements of HoneyBadgerBFT’s concrete implementation to showcase superior performance across varied settings.
Optimistic Fastlane Mechanisms. Certain works, such as those by Victor et al. [VICTOR01] and Cachin et al. [CACHIN02], have explored the incorporation of an optimistic ’fastlane’ mechanism, designed for operation within partially synchronous networks. However, in instances of sluggish agreement processes, these mechanisms transition to asynchronous Byzantine agreement protocols. Recent endeavors, such as those by Joleton [Joleton] and Boltdumbo [BOLTDUMBO], have adopted a similar strategy. Our work remains pertinent within this context as we rely on asynchronous protocols when the network exhibits Byzantine behavior, particularly within the pessimistic path.
Given our protocol’s focus on outputting requests from a single party, we omit comparisons with DAG-rider [DAG] and Aleph [ALEPH], which utilize directed acyclic graphs for agreement alongside sequential ACS. Additionally, recent efforts have aimed to eliminate the private setup phase in asynchronous Byzantine agreement protocols, emphasizing asynchronous distributed key generation [DIS01, DIS02, DIS03].
@@ 3 System Model
We use the standard notion [SECURE02, CACHIN01] to describe a distributed algorithm involving $n$ parties and an adversary in an authenticated setting. This section explains the foundational assumptions, network model, adversary model, computational model, and cryptographic tools used in our protocol.
@@ 3.1 System and Network Assumptions
Parties and Their Setup: There are $n$ parties, denoted as $p_{1},p_{2},\ldots,p_{n}$. We consider a trusted setup of threshold cryptosystems. The system provides a secret key and a public key to each party before the protocol starts. A party uses its secret key to sign a message, and other parties use the corresponding public key to verify the signed message. The generation and distribution of these keys are out of the scope of this paper. We follow standard literature for key generation and distribution and refer interested readers to [BORN01, BOLD01, Kate01]. We use the node and party alternatively throughout the paper.
@@
Computational Model: We adopt standard modern cryptographic assumptions and definitions from [SECURE02, CACHIN01]. The assumptions allow the parties and the adversary to be probabilistic polynomial-time interactive Turing machines. This means that upon receiving a message, a party carries out some computations, changes its state, generates outgoing messages if required, and waits for the next incoming message. To rule out infinite protocol executions and restrict the run time of the adversary, we require the message bits generated by honest parties to be probabilistic uniformly bounded by a polynomial in the security parameter $\lambda$. Therefore, we assume that the number of parties $n$ is bounded by a polynomial in $\lambda$.
@@ 3.2 Validated Asynchronous Byzantine Agreement:
Multi-valued Byzantine Agreement (MVBA) allows parties to agree on an arbitrary string $\{0,1\}^{l}$, which must satisfy a predefined validity condition before the honest parties agree on it. This protocol ensures that the agreed-upon values are valid even when inputs come from malicious parties. This property is crucial for ensuring the protocol’s correctness. Therefore, it is called the validated asynchronous byzantine agreement [CACHIN01, VABA, BYZ20].
@@ 3.3 Preliminaries
Guo et al. [FASTERDUMBO] constructed the $(1,\kappa,\epsilon)$-Committee Selection protocol using a threshold coin-tossing mechanism (see Appendix A.3), which is derived from threshold signatures. The protocol ensures that at least one honest party is a committee member with overwhelming probability $1-\epsilon-\text{neg}(\lambda)$, where $\text{neg}(\lambda)$ is a negligible function in the cryptographic security parameter $\lambda$, and $\epsilon$ is $exp(-\Omega\kappa)$.
Cryptographic Abstractions: We design a distributed algorithm in authenticated settings where we use robust, non-interactive threshold signatures to authenticate messages and a threshold coin-tossing protocol to select parties randomly [SECURE02].
@@
    Threshold Signature Scheme: We utilize a threshold signature scheme introduced in [THRESH01, SECURE02]. The basic idea is that there are $n$ parties, up to $f$ of which may be corrupted. The parties hold shares of the secret key of a signature scheme and may generate shares of signatures on individual messages. $t$ signature shares are both necessary and sufficient to construct a signature where $f<t\leq(n-f)$. The threshold signature scheme also provides a public key $pk$ along with secret key shares $sk_{1},\ldots,sk_{n}$, a global verification key $vk$, and local verification keys $vk_{1},\ldots,vk_{n}$. Initially, a party $p_{i}$ has information on the public key $vk$, a secret key $sk_{i}$, and the verification keys for all the parties’ secret keys. We describe the security properties of the scheme and related algorithms in Appendix A.2.
@@
    Threshold Coin-Tossing Scheme: The threshold coin-tossing scheme, as introduced in [THRESH01, SECURE02], involves parties holding shares of a pseudorandom function $F$ that maps the name $C$ (an arbitrary bit string) of a coin. We use a distributed pseudorandom function as a coin that produces $k^{\prime\prime}$ random bits simultaneously. The name $C$ is necessary and sufficient to construct the value $F(C)\in\{0,1\}^{k^{\prime\prime}}$ of the particular coin. The parties may generate shares of a coin — $t$ coin shares are both necessary and sufficient where $f<t\leq n-f$, similar to threshold signatures. The generation and verification of coin-shares are also non-interactive. We describe the security properties of the scheme and related algorithms in Appendix A.3.
4 Design of pMVBA
4.1 pMVBA Overview